Fermi National Laboratory


Accelerator Update

Monday January 12
The day shift began with store #3148 colliding, stacking in the Antiproton Source (Pbar), and with beam to MiniBoone.

Store # 3148 was terminated at 8:32 AM. Tevatron (Tev) studies are beginning.

At 9:52 AM cryo techs began replacing the D3 wet engine flywheel, the flywheel replacement was complete by 11:37 AM.

The Tev went into shot setup at 6:44 PM.

At 9:16 PM while initiating collisions the Tev quenched at A4, B1 and F4.

Pbar returned to stacking and quench investigation is underway.

At 11:41 PM the Tev is ramping once again.

Tuesday January 13
At 2:05 AM the RAW water system tripped off the MiniBoone horn. Miniboone will remain off to begin a 4 hour cooldown period followed by an access into M12 in the morning.

At 2:52 AM Booster power supply B:MH2 quit outputting current. Experts were called in to replace the SCR firing unit. The supply was back on by 4:54 AM.

At 7:47 AM water was added to the MiniBoone RAW system and MiniBoone began taking beam once again.

Tev quench investigation and Tev studies continue.

At 9:40 AM the MiniBoone experiment began experiencing problems with their data taking system. Beam was turned off while problem is being worked on.

At 12:00 PM an access was made into M12 to install 2 dehumidifiers.

MiniBoone turns back on at 2:39 PM.

At 3:49 PM the Tev studiers found the cause of the last Tev quench. The culprit was a vertical separator that was found to have the wrong polarity.

Shot setup began at 4:53 PM.

At 7:18 PM Operations established store #3157 with an initial luminosity of 17.14 E30.

Pbar then returned to stacking.

Pbar began shots to the Recycler Ring (RR) at 8:05 PM, shots were completed at 9:05 PM.

Pbar then returned to stacking.

Wednesday January 14
The Midnight shift began with store #3157 colliding, stacking, RR studies, and with beam to MiniBoone.

Store # 3157 was terminated at 6:01 AM.

Shot setup to the Tev began at 6:53 AM.

Today's Plans
Establish a Tev store, then return to stacking.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/14/2004   email Fermilab